40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Oct 20 11:31:06 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever 
    Reply  Tue Oct 26 10:01:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever TiltedCantilever.pdfcantileversurface.pdfclampsurface.pdf
       Reply  Thu Nov 18 11:09:24 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever CavitySetup.pdfGeometry01.jpeg
          Reply  Wed Dec 1 13:21:54 2021, shruti, Update, PSOMA, Towards locking PSOMA with cantilever IMG_1250.pngIMG_1252.pngIMG_1251.pngIMG_1253.pngIMG_1254.png
             Reply  Thu Dec 2 13:40:57 2021, shruti, Update, PSOMA, Progress and cavity parameters/ lock parameters 
Message ID: 2824     Entry time: Wed Oct 20 11:31:06 2021     Reply to this: 2825
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Towards locking PSOMA with cantilever 

[aaron, shruti]

Aaron removed the plastic wrap and began powering and reconnecting all electronics.

I drilled three holes into the plastic box that covers the cavity, reconnected the south laser diode and TEC to the ITC 502 combi controller and set the laser temperature to 9.888 kOhms and current to 126 mA (until the beam was visible).

 

21-Oct-21

Continued with alignment of the cavity until I observed a forest of peaks again while sinuoisoidally oscillating the temperature, although they can be observed even without temperature cycling now that we ahve a cantilever.

 

Next steps:

  • Mode-match to TEM00 better to be able to observe something on the IR camera
  • Drill holes for both the transmission PD (behind cantilever) and transmission monitor (behind curved mirror)
  • Set up locking electronics once again
  • Weigh and clamp down the plastic box
  • Procure new mount for the cube PBS to mount it in path so s-polarization transmits
  • Add PBS and HWP to the input path
  • Set up beat with another free-running laser to get a sense of the laser noise when measuring the noise spectra
ELOG V3.1.3-